# #EpiTwitter: Public Health Messaging During the COVID-19 Pandemic

### v1

@@ Abstract
Keywords Public health experts  $\cdot$ Public health messaging  $\cdot$ Pandemic  $\cdot$ Moral and emotional language
@@ Introduction
The COVID-19 pandemic created a world-wide public health crisis, disrupting daily lives and overwhelming healthcare facilities. During this time the need for communicating reliable medical information and public health guidance became very important. Social media platforms such as Twitter (now X) provided a space for public health experts (PHEs) in government, academia, and think tanks to communicate timely, reliable information about the pandemic to the public, to discuss research, and to provide guidance Tang et al. (2021); Abbas et al. (2021).
@@
However, as the pandemic progressed, discussions around the pandemic grew highly contentious and ideologically polarized Jiang et al. (2020); Rao et al. (2023). As public’s trust in institutions and experts eroded, health-related misinformation proliferated about all aspects of the pandemic, from its origins to alternative treatments and the efficacy of nonpharmaceutical interventions, and eventually the vaccine Rao et al. (2022). At the heart of this proliferation were influential pseudo-experts, like the “Disinformation Dozen” CCDH (2020), who amplified contrarian perspectives and challenged the recommendations made by PHEs. When PHEs advocated masking and social distancing, pseudo-experts promoted alternative treatments like Hydroxychloroquine and Ivermectin. When PHEs advocated vaccination as a method for reaching herd immunity, pseudo-experts emphasized vaccine injury.
@@
In this paper, we examine the messages shared by influential accounts posting about public health through the lens of affect. In psychology, “affect” is the experience of feeling or emotion, and it significantly shapes individual’s attitudes, beliefs and behaviors. In online interactions, affect influences how a message is crafted and how it resonates with audiences, ultimately affecting the message’s spread and impact. Research shows that people respond to the emotions expressed in online messages Kramer et al. (2014), although due to an asymmetry in human cognition Baumeister et al. (2001), posts expressing negative emotions receive more engagement than positive posts Coviello et al. (2014); Ferrara and Yang (2015). It has also been shown that emotionally charged messages, especially ones tapping into moral sentiments like outrage, spread farther online Brady et al. (2017, 2021). Affect provides reliable indicators for gauging public response to major events and policy decisions Renström and Bäck (2021); Sadler et al. (2005); Grover et al. (2019); Sterling and Jost (2018); Sagi and Dehghani (2014) and interacts with ideology to fuel polarization. In fact, political scientists have identified affective polarization—a phenomenon in which each party likes and trusts members of its own party by dislikes and distrusts members of the other party—as a fundamental threat to effective governance Iyengar and Westwood (2015); Iyengar et al. (2019). Public’s reactions to the pandemic, as measured via attitudes and sentiments expressed in online messages, were multifaceted Tsao et al. (2021) and grew polarized early in the pandemic Jiang et al. (2020). Moreover, there was an ideological asymmetry wherein conservatives shared more low quality health information than liberals Rao et al. (2021) and were also exposed to more misinformation Rao et al. (2022). In addition, conservatives expressed more negative moral sentiments in online posts about the pandemic than liberals Rao et al. (2023). However, to the best of our knowledge, few works have focused on online influencers and experts who shaped public health policy and disseminated health-related information to the public. As a result, we know little about the messaging strategies they used, the role that affect played in these messages, and how the public responded to the messages.
@@
We leverage methods introduced in Rao et al. (2023) to identify tweets about seven important pandemic-related issues: origins of the virus, lockdowna and stay-at-home orders, masking mandates, online schooling and education, healthcare, alternative treatments and therapeutics, and vaccines. We use state-of-the-art classifiers Alhuzali and Ananiadou (2021); Guo et al. (2023) to identify emotional and moral language used tweets. We then use regression to compare how affect shapes the health-related messages on different issues posted by PHEs and pseudo experts. Finally, we collect all replies for a sample of PHE tweets to study how the use of emotional and moral language impacts engagement by the public with these messages. These analyses yield the following key findings:
@@
Previous literature Slothuus and Bisgaard (2021); Bullock (2020) shows that the public follows cues from in-group elites and often oppose cues from the out-group. Effective messaging strategies can be crucial in times of public health crises. Individuals with higher COVID-19 knowledge practiced more protective behaviors Milich et al. (2024); Kim et al. (2021). Messaging that highlights risks to younger adults, in addition to risks to older adults was found to bring about a higher threat perception about COVID-19 Utych et al. (2020). On the other hand, messaging that appeals to the audience’s morals or fears for compliance was found to be polarising, divisive and undermined social cohesion McClaughlin et al. (2023). Latino vaccination rates saw significant increases in one Maryland County following the adoption of a cartoon grandmother WaPo (2021). Messaging comprised of vaccine safety and/or efficacy information and political leaders’ endorsement of vaccination were found to be highly effective strategies Xia and Nan (2023). Unvaccinated Pink et al. (2021) found that Republicans who were exposed to enodrsements from Republican elites witnessed higher vaccination intentions than those who viewed the Democratic elite endorsement with out-group elite exposures proving counter-productive.
@@
During the COVID-19 pandemic, social media was used by a large number of PHEs and healthcare providers to communicate best practices and information to the people, as well as to gain quicker access to healthcare information Goel and Gupta (2020). The pandemic, however, was a multifaceted public health crisis, presenting society with unprecedented challenges for which there was no established playbook. This coupled with the polarization Rao et al. (2021, 2022); Jiang et al. (2021) of the pandemic laid bare a fractured public health messaging apparatus Box-Steffensmeier and Moses (2021); Green et al. (2020); Van Scoy et al. (2021); Noar and Austin (2020). The emergence of contradictory theories and two polarized groups of influential elites and experts Mabrey (2021); Nogara et al. (2022), conspiracy theories about the origins of the pandemic, its severity and, the efficacy of prophylactic measures started to take hold Douglas (2021); Bavel et al. (2020). Initial theories revolved around the severity of the virus with several calling it a “hoax” and “plandemic” Frenkel et al. (2020). A study by the Pew Research Center Schaeffer (2020a) found that nearly 25% of the survey responders believed in the conspiracy theory propagated online that COVID-19 was probably created intentionally by powerful people. Another study found that nearly 3 in 10 Americans believe that COVID-19 was artificially created in a lab Schaeffer (2020b). Theories about virus transmission being connected to 5G, bats, pangolins and wet-markets were widely propagated by conspiracy theorists on social media Langguth et al. (2023); Taylor (2020). As the pandemic progressed, we also witnessed the propagation of pseudo-scientific cures for COVID-19 Robins-Early (2020); Woo (2021). With increased pandemic related engagement from the general public, these conspiracy theories soon started to proliferate on social media platforms Desai et al. (2022); Kearney et al. (2020); Ahmed et al. (2020); Gruzd and Mai (2020); Ferrara (2020); Gottlieb and Dyer (2020). Antonakis (2021) highlights the role of influential accounts in mitigation efforts. As influential elites, often holding advanced medical degrees began contradicting with other PHEs on various aspects of the pandemic, consensus amongst the general public suffered as a consequence and at times resulted in grave outcomes Geleris et al. (2020); Sehgal et al. (2022); Pradelle et al. (2024).
@@
Expression of fear and anger were found to indicate support for restrictive COVID-19 mitigation policies such as lockdowns to limit the spread of COVID-19 Renström and Bäck (2021), while anxiety predicted support for economic policies. Anger was found to indicate support for aggressive responses to transgressors Sadler et al. (2005); Skitka et al. (2006). Hatemi et al. (2013) found fear to be a strong underlying factor in anti-immigration and pro-segregation stances. Previous studies APA (2021) relied on surveys to show an increase in distress and uncertainty during the pandemic. Agrawal et al. (2022) investigated sentiments of Tweets about the COVID-19 vaccine, post-Covid health factors, and health service providers. Among the three topics, healthcare providers had the largest positive sentiment, resulting in an inference that posters were happy with their care and appreciated the work of health providers. Lwin et al. (2020) found that public emotions in Twitter shifted from fear to anger early in the pandemic. Wheaton et al. (2021) revealed that greater susceptibility to emotion contagion was associated with greater concern about the spread of COVID-19. The care and fairness moral foundations were found to correlate with compliance of COVID-19 health recommendations including masking, staying at home and social distancing Díaz and Cova (2022); Chan (2021). Moral attitudes were also able to predict county-level vaccination rate Reimer et al. (2022) and vaccine hesitancy Nan et al. (2022). Pacheco et al. (2022) found that care/harm was associated with pro-vaccine sentiment, whereas liberty/oppression was correlated with anti-vaccine attitudes. While vaccinations are a critical polarizing issue in the discussion of COVID-19, no work as of yet has explored differences in moral appeals across a broader range of contentious COVID-19 issues.
@@
Social sharing of opinions and emotions is ubiquitous, and social media has greatly expanded its scope John (2013); Bazarova et al. (2015). Bazarova et al. (2015) investigated how responses to what a user shared affected their feeling of satisfaction. The authors argue that while sharing emotions with large groups of users online is often viewed as over-sharing, it is likely a means to increase the likelihood of receiving responses from virtual listeners. Analyzing Facebook status updates, Burke and Develin (2016) found that posts with positive emotions received more likes. The comments associated with these posts were also more positive Burke and Develin (2016). Positive emotion words were also shown to have a positive correlation with the number of retweets Kim and Yoo (2012). Sousa et al. (2010) reports that while social connections dominate reply behavior, for authors with large ego networks, there is a separation between who replies based on the topic of the post. Early thematic analysis of public replies to the pandemic found themes of prevention, symptoms, views on politicians and, humor Bojja et al. (2020). Replies by anti-vaxx users were found to be more toxic than users with other beliefs about vaccines Miyazaki et al. (2021). Gallagher et al. (2021) found groups to preferentially amplify elites that are demographically similar to them.
@@
We identified a seed set of 30 PHEs and 30 pseudo-experts who were active on social media platforms during the pandemic (see Table 1. While the seed set of PHEs comprised largely of individuals who hold advanced degrees in medicine, epidemiology, genomics, infectious diseases or vaccine development, allowing them to offer well-informed, scientifically grounded perspectives on the pandemic, pseudo-experts include individuals with or without medical degrees, proposing pseudo-scientific theories, alternative treatments and therapeutics. The group also includes the “Disinformation Dozen”, a group of individuals and organizations identified by the Center for Countering Digital Hate as being responsible for promoting false claims about COVID-19 CCDH (2020).
@@
Table 1: Twitter Handles of Accounts Associated with Public Health Experts and Pseudo Experts
Figure 1: Retweet network. Nodes represent PHEs (green) and pseudo-experts (orange) and retweet interactions between them. Green edges represent interactions where a PHE was retweeted, and orange edges represent interactions where a pseudo-expert was retweeted. The size of the node is proportional to the number of times the expert was retweeted.
In order to expand these sets of experts, we rely on a publicly available COVID-19 Twitter dataset Chen et al. (2020) comprising of over 1B COVID-19 tweets gathered between January 21, 2020 to January 20, 2021. Previous studies have shown that retweet interactions where users repost or reshare content originally generated by other users, indicates an endorsement of the content Metaxas et al. (2015); Boyd et al. (2010). Working with the assumption that individuals often agree with others who share identical beliefs Rao et al. (2022); Cinelli et al. (2021); Colleoni et al. (2014); Kang and Lerman (2012), we isolate retweet interactions involving PHEs and pseudo-experts to construct two networks — one involving accounts retweeted by PHEs and the other accounts retweeted by pseudo-experts. We use Eigenvector centrality Ghosh and Lerman (2010) to rank influential accounts who are retweeted by at least one PHE or pseudo-expert, and select the top-500 most influential accounts. We remove organizations from this list, which leaves us with $489$ and $356$ accounts corresponding to individual PHEs and pseudo-experts respectively. Having identified the group of health and pseudo-experts, we then extract their tweets. Overall, we have $340K$ tweets from PHEs and $175K$ tweets from pseudo-experts.
@@
Further analyzing the content shared by PHEs and pseudo experts, we extract the URLs they shared in their posts and identify the pay-level domains (PLDs) these URLs point to. We compute the log-odds ratio to identify which group is more likely to share each PLD. Supplementary Fig. 12(a) shows the top-15 PLDs for both groups. we find that PHEs were more likely to share URLs from highly reputable sources such as Journal of American Medical Association, Nature, Boston Review and New York Times. On the other hand, pseudo experts share more questionable sources such as The Gateway Pundit, Childrens Health Defense, Patriot Project and Russia Today among others. PLDs often have ideological leanings, ranging from liberal (00) to least-biased ($0.5$) to conservative ($1$) Check (2023). Supplementary Fig. 12(b) compares the distribution of ideological leanings of information sources shared by PHEs and pseudo-experts. While, PHEs tended to share more liberal sources, pseudo experts shared more conservative sources.
@@
We decompose the multi-faceted discussion about the COVID-19 pandemic along various contentious issues: COVID-19 origins, lockdowns and business closures, mask mandates, school closures, therapeutics, healthcare and, vaccines. In order to do so, we rely on methods discussed in Rao et al. (2023); Eisenstein et al. (2011) to extract issue-relevant keywords from Wikipedia articles. Once we identify keywords, we identify tweets that explicitly mention any of these keywords as being issue-relevant. This approach was validated in Rao et al. (2023) as being able to accurately identify issue-relevant content. Table 2 illustrates sample tweets from our dataset discussing each issue.
@@
Table 2: Sample tweets for each COVID-19 issue.
@@
In order to identify emotions expressed in tweets and replies, we employed SpanEmo  Alhuzali and Ananiadou (2021), a state-of-the-art multi-label emotion detection model. This model was fine-tuned using the SemEval 2018 Task 1e-c dataset Mohammad et al. (2018). SpanEmo, using a transformer-based architecture. It surpasses previous methods in its ability to capture the correlations among various emotions. When presented with the text of a tweet, the model generates confidence scores for the presence of a wide spectrum of emotions. We later bin these confidence scores using a $0.5$ threshold to binarize the output. The emotions it can identify include anticipation, joy, love, optimism, anger, disgust, fear, sadness and, pessimism. The definitions for these emotions are borrowed from Mohammad et al. (2018) as follows:
@@
Prior research has shown that emotional and moral language in social media messages impacts how they are received by the audiences and engagement Brady et al. (2021); Kramer et al. (2014). The Moral Foundations Theory (MFT)  Haidt and Graham (2007) provides a framework for understanding how moral values shape people’s political attitudes and behaviors. MFT proposes that individuals’ values and judgments can be described by five moral virtue/vice pairs - care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation. More specifically,
@@
Our morality detection model is trained on the BERT transformer-based pretrained language model by Devlin et al. (2018). The training process involves three Twitter datasets, a manually annotated COVID dataset Rojecki et al. (2021), the Moral Foundation Twitter Corpus dataset covering six different topics Hoover et al. (2020), and a dataset of political tweets from US congress members Johnson and Goldwasser (2018). By incorporating an in-domain training set focused on COVID-19, along with other diverse datasets spanning various topics, we enhance the model’s generalizability for application to target data as discussed in Guo et al. (2023).

### v2

@@ Abstract
keywords:
Public health experts Public health messaging Pandemic Moral and emotional language
@@ Introduction
The COVID-19 pandemic created a world-wide public health crisis, disrupting daily lives and overwhelming healthcare facilities. During this time the need for communicating reliable medical information and public health guidance became very important. Social media platforms such as Twitter (now X) provided a space for public health experts (PHEs) in government, academia, and think tanks to communicate timely, reliable information about the pandemic to the public, to discuss research, and to provide guidance [1, 2].
@@
However, as the pandemic progressed, discussions around the pandemic grew highly contentious and ideologically polarized [3, 4]. As public’s trust in institutions and experts eroded, health-related misinformation proliferated about all aspects of the pandemic, from its origins to alternative treatments and the efficacy of nonpharmaceutical interventions, and eventually the vaccine [5]. At the heart of this proliferation were influential pseudo-experts, like the “Disinformation Dozen” [6], who amplified contrarian perspectives and challenged the recommendations made by PHEs. When PHEs advocated masking and social distancing, pseudo-experts promoted alternative treatments like Hydroxychloroquine and Ivermectin. When PHEs advocated vaccination as a method for reaching herd immunity, pseudo-experts emphasized vaccine injury.
@@
In this paper, we examine the messages shared by influential accounts posting about public health through the lens of affect. In psychology, “affect” is the experience of feeling or emotion, and it significantly shapes individual’s attitudes, beliefs and behaviors. In online interactions, affect influences how a message is crafted and how it resonates with audiences, ultimately affecting the message’s spread and impact. Research shows that people respond to the emotions expressed in online messages [7], although due to an asymmetry in human cognition [8], posts expressing negative emotions receive more engagement than positive posts [9, 10]. It has also been shown that emotionally charged messages, especially ones tapping into moral sentiments like outrage, spread farther online [11, 12]. Affect provides reliable indicators for gauging public response to major events and policy decisions [13, 14, 15, 16, 17] and interacts with ideology to fuel polarization. In fact, political scientists have identified affective polarization—a phenomenon in which each party likes and trusts members of its own party by dislikes and distrusts members of the other party—as a fundamental threat to effective governance [18, 19]. Public’s reactions to the pandemic, as measured via attitudes and sentiments expressed in online messages, were multifaceted [20] and grew polarized early in the pandemic [3]. Moreover, there was an ideological asymmetry wherein conservatives shared more low quality health information than liberals [21] and were also exposed to more misinformation [5]. In addition, conservatives expressed more negative moral sentiments in online posts about the pandemic than liberals [4]. However, to the best of our knowledge, few works have focused on online influencers and experts who shaped public health policy and disseminated health-related information to the public. As a result, we know little about the messaging strategies they used, the role that affect played in these messages, and how the public responded to the messages.
@@
We leverage methods introduced in [4] to identify tweets about seven important pandemic-related issues: origins of the virus, lockdowna and stay-at-home orders, masking mandates, online schooling and education, healthcare, alternative treatments and therapeutics, and vaccines. We use state-of-the-art classifiers [22, 23] to identify emotional and moral language used tweets. We then use regression to compare how affect shapes the health-related messages on different issues posted by PHEs and pseudo experts. Finally, we collect all replies for a sample of PHE tweets to study how the use of emotional and moral language impacts engagement by the public with these messages. These analyses yield the following key findings:
@@
Previous literature [24, 25] shows that the public follows cues from in-group elites and often oppose cues from the out-group. Effective messaging strategies can be crucial in times of public health crises. Individuals with higher COVID-19 knowledge practiced more protective behaviors [26, 27]. Messaging that highlights risks to younger adults, in addition to risks to older adults was found to bring about a higher threat perception about COVID-19 [28]. On the other hand, messaging that appeals to the audience’s morals or fears for compliance was found to be polarising, divisive and undermined social cohesion [29]. Latino vaccination rates saw significant increases in one Maryland County following the adoption of a cartoon grandmother [30]. Messaging comprised of vaccine safety and/or efficacy information and political leaders’ endorsement of vaccination were found to be highly effective strategies [31]. Unvaccinated [32] found that Republicans who were exposed to enodrsements from Republican elites witnessed higher vaccination intentions than those who viewed the Democratic elite endorsement with out-group elite exposures proving counter-productive.
@@
During the COVID-19 pandemic, social media was used by a large number of PHEs and healthcare providers to communicate best practices and information to the people, as well as to gain quicker access to healthcare information [33]. The pandemic, however, was a multifaceted public health crisis, presenting society with unprecedented challenges for which there was no established playbook. This coupled with the polarization [21, 5, 34] of the pandemic laid bare a fractured public health messaging apparatus [35, 36, 37, 38]. The emergence of contradictory theories and two polarized groups of influential elites and experts [39, 40], conspiracy theories about the origins of the pandemic, its severity and, the efficacy of prophylactic measures started to take hold [41, 42]. Initial theories revolved around the severity of the virus with several calling it a “hoax” and “plandemic” [43]. A study by the Pew Research Center [44] found that nearly 25% of the survey responders believed in the conspiracy theory propagated online that COVID-19 was probably created intentionally by powerful people. Another study found that nearly 3 in 10 Americans believe that COVID-19 was artificially created in a lab [45]. Theories about virus transmission being connected to 5G, bats, pangolins and wet-markets were widely propagated by conspiracy theorists on social media [46, 47]. As the pandemic progressed, we also witnessed the propagation of pseudo-scientific cures for COVID-19 [48, 49]. With increased pandemic related engagement from the general public, these conspiracy theories soon started to proliferate on social media platforms [50, 51, 52, 53, 54, 55]. [56] highlights the role of influential accounts in mitigation efforts. As influential elites, often holding advanced medical degrees began contradicting with other PHEs on various aspects of the pandemic, consensus amongst the general public suffered as a consequence and at times resulted in grave outcomes [57, 58, 59].
@@
Expression of fear and anger were found to indicate support for restrictive COVID-19 mitigation policies such as lockdowns to limit the spread of COVID-19 [13], while anxiety predicted support for economic policies. Anger was found to indicate support for aggressive responses to transgressors [14, 60]. [61] found fear to be a strong underlying factor in anti-immigration and pro-segregation stances. Previous studies [62] relied on surveys to show an increase in distress and uncertainty during the pandemic. [63] investigated sentiments of Tweets about the COVID-19 vaccine, post-Covid health factors, and health service providers. Among the three topics, healthcare providers had the largest positive sentiment, resulting in an inference that posters were happy with their care and appreciated the work of health providers. [64] found that public emotions in Twitter shifted from fear to anger early in the pandemic. [65] revealed that greater susceptibility to emotion contagion was associated with greater concern about the spread of COVID-19. The care and fairness moral foundations were found to correlate with compliance of COVID-19 health recommendations including masking, staying at home and social distancing [66, 67]. Moral attitudes were also able to predict county-level vaccination rate [68] and vaccine hesitancy [69]. [70] found that care/harm was associated with pro-vaccine sentiment, whereas liberty/oppression was correlated with anti-vaccine attitudes. While vaccinations are a critical polarizing issue in the discussion of COVID-19, no work as of yet has explored differences in moral appeals across a broader range of contentious COVID-19 issues.
@@
Social sharing of opinions and emotions is ubiquitous, and social media has greatly expanded its scope [71, 72]. [72] investigated how responses to what a user shared affected their feeling of satisfaction. The authors argue that while sharing emotions with large groups of users online is often viewed as over-sharing, it is likely a means to increase the likelihood of receiving responses from virtual listeners. Analyzing Facebook status updates, [73] found that posts with positive emotions received more likes. The comments associated with these posts were also more positive [73]. Positive emotion words were also shown to have a positive correlation with the number of retweets [74]. [75] reports that while social connections dominate reply behavior, for authors with large ego networks, there is a separation between who replies based on the topic of the post. Early thematic analysis of public replies to the pandemic found themes of prevention, symptoms, views on politicians and, humor [76]. Replies by anti-vaxx users were found to be more toxic than users with other beliefs about vaccines [77]. [78] found groups to preferentially amplify elites that are demographically similar to them.
@@
We identified a seed set of 30 PHEs and 30 pseudo-experts who were active on social media platforms during the pandemic (see Table 1. While the seed set of PHEs comprised largely of individuals who hold advanced degrees in medicine, epidemiology, genomics, infectious diseases or vaccine development, allowing them to offer well-informed, scientifically grounded perspectives on the pandemic, pseudo-experts include individuals with or without medical degrees, proposing pseudo-scientific theories, alternative treatments and therapeutics. The group also includes the “Disinformation Dozen”, a group of individuals and organizations identified by the Center for Countering Digital Hate as being responsible for promoting false claims about COVID-19 [6].
@@
Table 1: Twitter Handles. Twitter Handles of Accounts Associated with Public Health Experts and Pseudo Experts
Figure 1: Retweet interactions. Nodes represent PHEs (green) and pseudo-experts (orange) and retweet interactions between them. Green edges represent interactions where a PHE was retweeted, and orange edges represent interactions where a pseudo-expert was retweeted. The size of the node is proportional to the number of times the expert was retweeted.
In order to expand these sets of experts, we rely on a publicly available COVID-19 Twitter dataset [79] comprising of over 1B COVID-19 tweets gathered between January 21, 2020 to January 20, 2021. Previous studies have shown that retweet interactions where users repost or reshare content originally generated by other users, indicates an endorsement of the content [80, 81]. Working with the assumption that individuals often agree with others who share identical beliefs [5, 82, 83, 84], we isolate retweet interactions involving PHEs and pseudo-experts to construct two networks — one involving accounts retweeted by PHEs and the other accounts retweeted by pseudo-experts. We use Eigenvector centrality [85] to rank influential accounts who are retweeted by at least one PHE or pseudo-expert, and select the top-500 most influential accounts. We remove organizations from this list, which leaves us with $489$ and $356$ accounts corresponding to individual PHEs and pseudo-experts respectively. Having identified the group of health and pseudo-experts, we then extract their tweets. Overall, we have $340K$ tweets from PHEs and $175K$ tweets from pseudo-experts.
@@
Further analyzing the content shared by PHEs and pseudo experts, we extract the URLs they shared in their posts and identify the pay-level domains (PLDs) these URLs point to. We compute the log-odds ratio to identify which group is more likely to share each PLD. Supplementary Figure 12(a) shows the top-15 PLDs for both groups. we find that PHEs were more likely to share URLs from highly reputable sources such as Journal of American Medical Association, Nature, Boston Review and New York Times. On the other hand, pseudo experts share more questionable sources such as The Gateway Pundit, Childrens Health Defense, Patriot Project and Russia Today among others. PLDs often have ideological leanings, ranging from liberal (00) to least-biased ($0.5$) to conservative ($1$) [86]. Supplementary Figure 12(b) compares the distribution of ideological leanings of information sources shared by PHEs and pseudo-experts. While, PHEs tended to share more liberal sources, pseudo experts shared more conservative sources.
@@
We decompose the multi-faceted discussion about the COVID-19 pandemic along various contentious issues: COVID-19 origins, lockdowns and business closures, mask mandates, school closures, therapeutics, healthcare and, vaccines. In order to do so, we rely on methods discussed in [4, 87] to extract issue-relevant keywords from Wikipedia articles. Once we identify keywords, we identify tweets that explicitly mention any of these keywords as being issue-relevant. This approach was validated in [4] as being able to accurately identify issue-relevant content. Table 2 illustrates sample tweets from our dataset discussing each issue.
@@
Table 2: Issue relevant tweets. Sample tweets for each COVID-19 issue.
@@
In order to identify emotions expressed in tweets and replies, we employed SpanEmo  [22], a state-of-the-art multi-label emotion detection model. This model was fine-tuned using the SemEval 2018 Task 1e-c dataset [88]. SpanEmo, using a transformer-based architecture. It surpasses previous methods in its ability to capture the correlations among various emotions. When presented with the text of a tweet, the model generates confidence scores for the presence of a wide spectrum of emotions. We later bin these confidence scores using a $0.5$ threshold to binarize the output. The emotions it can identify include anticipation, joy, love, optimism, anger, disgust, fear, sadness and, pessimism. The definitions for these emotions are borrowed from [88] as follows:
@@
Prior research has shown that emotional and moral language in social media messages impacts how they are received by the audiences and engagement [12, 7]. The Moral Foundations Theory (MFT)  [89] provides a framework for understanding how moral values shape people’s political attitudes and behaviors. MFT proposes that individuals’ values and judgments can be described by five moral virtue/vice pairs - care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation. More specifically,
@@
Our morality detection model is trained on the BERT transformer-based pretrained language model by [90]. The training process involves three Twitter datasets, a manually annotated COVID dataset [91], the Moral Foundation Twitter Corpus dataset covering six different topics [92], and a dataset of political tweets from US congress members [93]. By incorporating an in-domain training set focused on COVID-19, along with other diverse datasets spanning various topics, we enhance the model’s generalizability for application to target data as discussed in [23].